OpenStudy (anonymous):

The width of a rectangle is 18 in. less than its length. If the area of the rectangle is 1440 in^2, find the dimensions of the rectangle.

OpenStudy (anonymous):

30in x 48in = 1440in^2

Directrix (directrix):

If l is the length, then the width is (l - 18) Area of a rectangle = lw A = l ((l - 18)) 1440 = L^2 - 18L L^2 - 18L = 1440 L^2 - 18L - 1440 = 0 (L - 48) (L + 30) = 0 L = 48 or L = -30 L = 48 and L - 18 = 30 48 by 30

OpenStudy (anonymous):

additonal: we can't have -30 as the value for length for we're dealing with length and measure, there is no negative length in real application. .